# Spectral Editing of Activations for Large Language Model Alignment

### v1

@@ Abstract
Large language models (LLMs) often exhibit undesirable behaviours, such as generating untruthful or biased content. Editing their internal representations has been shown to be effective in mitigating such behaviours on top of the existing alignment methods. We propose a novel inference-time editing method, namely spectral editing of activations (SEA), to project the input representations into directions with maximal covariance with the positive demonstrations (e.g., truthful) while minimising covariance with the negative demonstrations (e.g., hallucinated). We also extend our method to non-linear editing using feature functions. We run extensive experiments on benchmarks concerning truthfulness and bias with six open-source LLMs of different sizes and model families. The results demonstrate the superiority of SEA in effectiveness, generalisation to similar tasks, as well as inference and data efficiency. We also show that SEA editing only has a limited negative impact on other model capabilities.  [^1] Our code and edited models are available on https://github.com/yfqiu-nlp/sea-llm.
@@ 3.1 Truthfulness
Datasets. We evaluate all compared methods on TruthfulQA [22], which consists of 817 questions in 38 subcategories, each paired with a single best answer, and multiple correct/incorrect answers. TruthfulQA contains two evaluation protocols: multiple-choice question answering, and generation. We mainly use the first to ensure the comparability with previous work [20, 21, 6]. We show an example of a TruthfulQA data instance together with the demonstrations we used in Appendix E.
@@ 3.2 Bias
Dataset. We assess the model bias on the Bias Benchmark for QA (BBQ; Parrish et al. 25), which is widely adopted in LLM evaluation [17, 29, 2]. BBQ formulates bias evaluation as a question-answering task, allowing us to easily construct the paired positive and negative demonstrations. BBQ contains 29,246 questions covering 11 diverse types of common bias. We randomly sampled 5,246 questions for evaluation, and the rest were used for training and validation purposes. We use the disambiguated version for a more comprehensive evaluation, providing the model with sufficiently informative contexts. This approach enables us to assess whether the model biases influence its selection of a correct answer choice. See Appendix E for example demonstrations.
@@ 4.1 Truthfulness Evaluation
Method MC1 MC2 Time $\times$ 100
ICL (LLaMA-2) 28.39 43.42 -
ICL (LLaMA-2-Chat) 36.96 54.68 4.90
LoRA-FT (LLaMA-2-Chat; $N=1000$) 35.74 54.61 5.16
LoRA-FT (LLaMA-2-Chat; $N=2000$) 35.01 54.24 5.04
ITI 37.01 54.66 5.82
DoLA 32.97 60.84 5.60
CD (13B-Chat vs. 7B-Chat) 28.15 54.87 -
ICD (Prompt-version) 37.87 57.77 9.67
SEA ($N=1000,K=99\%,L=4$) 38.31 55.27 5.08
SEA ($N=2000,K=99.8\%,L=21$) 39.41 57.15 5.93
Table 1: Results for truthfulness evaluation on TruthfulQA using multiple-choice-based metrics (MC1/2). All baselines are built on top of LLaMA-2-Chat-7B as an LLM backbone. We measure the inference efficiency for each method in average inference time (seconds) per sample. Part of the listed results are from [40].
The best result is from our proposed SEA method, which outperforms ICL by 2.45 and 2.47 points for MC1 and MC2, respectively. When compared to alternative approaches, SEA achieves the highest MC1 score while incurring the minimal sacrifice in inference speed when we only modify the last four layers. For instance, SEA’s modification results in only a 3.67% increase in inference speed, contrasting with significantly larger increases of 18.78%, 14.29%, and 97.34% for ITI, DoLA, and ICD, respectively.
@@ 4.2 Bias Evaluation
(c) Left: Accuracy of all methods on each group of bias type in BBQ. Right: results on the whole BBQ testing set from applying various methods on LLaMA-2-Chat-7B. For accuracy (A%), higher values are better; for unknown-answer response rate (U%), bias score (BS%) and stereotypical response rate (SR%), lower is better. We bold the best method in each column, and mark the methods that improve ICL positively.
@@ 4.4 Post-Editing Performance on Control Tasks
Table 4: Performance of LLaMA-2-Chat-7B and its three SEA edited models for truthfulness and fairness on six control tasks covering multi-task ability, commonsense question answering, and mathematical ability. Details of evaluation are provided in Appendix C.2.
@@ 5 Related Work
Modifying the activation of a trained model, thus altering the model’s behaviour [28, 15, 23] or internal knowledge [8], represents a lightweight method to exert control over the model’s generation. Li et al. [20] propose an approach to initially identify which attention heads of an LLM are accountable for generating hallucinations, followed by editing activations toward truthful directions during inference. Another way is extracting latent vectors directly from the trained model and leveraging these vectors to regulate the model’s inference [38, 34, 30]. Recently, Singh et al. [33] demonstrated the efficacy of fitting an optimal transport from negative to positive activations to facilitate effective non-linear editing. Activation editing finds application in decoding as well, either by contrasting activations from various layers [6] or by using a weaker model to edit activations from a stronger model [21, 40].
@@ Appendix B Qualitative Analysis
Appendix C Implementation Details
C.1 Main Experiments for LLaMA-2-Chat-7B Model
@@
Linear SEA for Fairness. For the best linear SEA for fairness, we use 1000 pairs of demonstrations randomly sampled from our training split of BBQ to obtain the editing projections for truthfulness. We set the hyperparameter $K=99.9\%$ and we edit the top 3 layers.
$\Phi-$SEA for Fairness. For the non-linear SEA for fairness using squared-exponential and hyperbolic tangent feature functions, we use 1000 pairs of demonstration randomly sampled from our training split of BBQ to obtain the editing projections for truthfulness. We set the hyperparameter $K=99.99\%$ and we edit the top 2 layers.
@@
C.2 Experiments Setup for Control Tasks
@@
Appendix D Limitations
@@
Appendix E Demonstrations for TruthfulQA and BBQ

### v2

@@ Abstract
Large language models (LLMs) often exhibit undesirable behaviours, such as generating untruthful or biased content. Editing their internal representations has been shown to be effective in mitigating such behaviours on top of the existing alignment methods. We propose a novel inference-time editing method, namely spectral editing of activations (SEA), to project the input representations into directions with maximal covariance with the positive demonstrations (e.g., truthful) while minimising covariance with the negative demonstrations (e.g., hallucinated). We also extend our method to non-linear editing using feature functions. We run extensive experiments on benchmarks concerning truthfulness and bias with six open-source LLMs of different sizes and model families. The results demonstrate the superiority of SEA in effectiveness, generalisation to similar tasks, as well as computation and data efficiency. We also show that SEA editing only has a limited negative impact on other model capabilities.  [^1] Our code and edited models are available on https://github.com/yfqiu-nlp/sea-llm .
@@ 3.1 Truthfulness
Datasets. We evaluate all compared methods on TruthfulQA [22], which consists of 817 questions in 38 subcategories, each paired with a single best answer, and multiple correct/incorrect answers. TruthfulQA contains two evaluation protocols: multiple-choice question answering, and generation. We mainly use the first to ensure the comparability with previous work [20, 21, 6]. We show an example of a TruthfulQA data instance together with the demonstrations we used in Appendix H.
@@ 3.2 Bias
Dataset. We assess the model bias on the Bias Benchmark for QA (BBQ; Parrish et al. 25), which is widely adopted in LLM evaluation [17, 29, 2]. BBQ formulates bias evaluation as a question-answering task, allowing us to easily construct the paired positive and negative demonstrations. BBQ contains 29,246 questions covering 11 diverse types of common bias. We randomly sampled 5,246 questions for evaluation, and the rest were used for training and validation purposes. We use the disambiguated version for a more comprehensive evaluation, providing the model with sufficiently informative contexts. This approach enables us to assess whether the model biases influence its selection of a correct answer choice. See Appendix H for example demonstrations.
@@ 4.1 Truthfulness Evaluation
[…]
Method MC1 MC2 TInf. TTrain
ICL (LLaMA-2) 28.39 43.42 - -
ICL (LLaMA-2-Chat) 36.96 54.68 4.90 -
LoRA-FT (LLaMA-2-Chat; $N=1000$) 35.74 54.61 5.16 299
LoRA-FT (LLaMA-2-Chat; $N=2000$) 35.01 54.24 5.04 1190
ITI 37.01 54.66 5.82 -
DoLA 32.97 60.84 5.60 -
CD (13B-Chat vs. 7B-Chat) 28.15 54.87 - -
ICD (Prompt-version) 37.87 57.77 9.67 -
SEA ($N=1000,K=99\%,L=4$) 38.31 55.27 5.08 140
SEA ($N=2000,K=99.8\%,L=21$) 39.41 57.15 5.93 152
The best result is from our proposed SEA method, which outperforms ICL by 2.45 and 2.47 points for MC1 and MC2, respectively. When compared to alternative approaches, SEA achieves the highest MC1 score while incurring the minimal sacrifice in inference speed when we only modify the last four layers. SEA has only a 3.67% increase in inference speed, contrasting with the larger increases of 18.78%, 14.29%, and 97.34% for ITI, DoLA, and ICD, respectively. We also highlight the training efficiency, i.e., computing editing projections, of SEA compared to gradient-based optimisation methods (e.g., LoRA), an advantage that becomes more significant with an increasing number of demonstrations.
@@ 4.2 Bias Evaluation
[…]
@@ 4.4 Post-Editing Performance on Control Tasks
Table 4: Performance of LLaMA-2-Chat-7B and its three SEA edited models for truthfulness and fairness on six control tasks covering multi-task ability, commonsense question answering, and mathematical ability. Details of evaluation are provided in Appendix D.4.
@@ 5 Related Work
Modifying the activation of a trained model, thus altering the model’s behaviour [28, 15, 23, 43] or internal knowledge [8], represents a lightweight method to control the model’s generation. Li et al. [20] propose an approach to initially identify which attention heads of an LLM are accountable for generating hallucinations, followed by editing activations toward truthful directions during inference. Another way is extracting latent vectors directly from the trained model and leveraging these vectors to regulate the model’s inference [38, 34, 30]. Recently, Singh et al. [33] demonstrated the efficacy of fitting an optimal transport from negative to positive activations to facilitate effective non-linear editing. Activation editing finds application in decoding as well, either by contrasting activations from various layers [6] or by using a weaker model to edit activations from a stronger model [21, 40].
@@ Appendix B Qualitative Analysis
Appendix C Analysis in the Effect of $K$
Figure 6: Analysis in the effect of hyperparameter $K$. Values in the y-axis represent the model performance in MC1 (top panel) and MC2 (bottom panel). $K$ values are on the x-axis, chosen form {95%, 99%, 99.5%, 99.99%}. We also alter the number of editing layers ($L$) to verify if the observation can be generalised to different settings. The used LLM in this analysis is the LLaMA-2-Chat-7B.
In this section, we explore the impact of hyperparameter $K$ on the performance of the SEA-edited LLaMA-2-Chat-7B model (MC-1/2) on TruthfulQA. Specifically, we vary $K$ across values of $\{95\%,99\%,99.5\%,99.99\%\}$, alongside $L$ values from $\{1,2,3,4,5\}$, and present the corresponding MC1 and MC2 scores for each experimental configuration.
Our analysis reveals a consistent pattern: as $K$ increases, the model’s performance initially improves, reaching its peak around 99.5%, before declining thereafter. We interpret this trend as follows: below a certain threshold, increasing $K$ allows SEA’s projections to capture more information related to positive demonstrations and less information related to negative demonstrations. However, as $K$ surpasses the turning point, the heightened emphasis on positive signals leads to the incorporation of noise in the positive demonstrations, also diminishing performance by reducing task-related information from negative demonstrations.
Appendix D Implementation Details
D.1 Main Experiments for LLaMA-2-Chat-7B Model
@@
Linear SEA for Fairness. For the best linear SEA for fairness, we use 1000 pairs of demonstrations randomly sampled from our training split of BBQ to obtain the editing projections for fairness. We set the hyperparameter $K=99.9\%$ and we edit the top 3 layers.
$\Phi-$SEA for Fairness. For the non-linear SEA for fairness using squared-exponential and hyperbolic tangent feature functions, we use 1000 pairs of demonstration randomly sampled from our training split of BBQ to obtain the editing projections for fairness. We set the hyperparameter $K=99.99\%$ and we edit the top 2 layers.
@@
D.2 Editing Fairness for Other LLMs
LLaMA-2-Chat 13B Model. For the best linear SEA setting, we use 1000 pairs of demonstrations randomly sampled from our training split of BBQ to obtain the editing projections for fairness. We set the hyperparameter $K=99.9\%$ and we edit the top 1 layer. For the best $\Phi-$SEA setting, we use squared-exponential feature function, $K=99.99\%$, edit the top 2 layers, and use 1000 pairs of demonstrations.
LLaMA-2-Chat 70B Model. For the best linear SEA setting, we use 1000 pairs of demonstrations randomly sampled from our training split of BBQ to obtain the editing projections for fairness. We set the hyperparameter $K=99.9\%$ and we edit the top 1 layer. For the best $\Phi-$SEA setting, we use hyperbolic tangent feature function, $K=99.9\%$, edit the top 1 layer, and use 1000 pairs of demonstrations.
Gemma-it 2B Model. For the best linear SEA setting, we use 1000 pairs of demonstrations randomly sampled from our training split of BBQ to obtain the editing projections for fairness. We set the hyperparameter $K=99.9\%$ and we edit the top 3 layers. For the best $\Phi-$SEA setting, we use ELU feature function, $K=99.99\%$, edit the top 1 layer, and use 1000 pairs of demonstrations.
Gemma-it 7B Model. For the best linear SEA setting, we use 1000 pairs of demonstrations randomly sampled from our training split of BBQ to obtain the editing projections for fairness. We set the hyperparameter $K=99.9\%$ and we edit the top 2 layers. For the best $\Phi-$SEA setting, we use squared-exponential feature function, $K=99.99\%$, edit the top 3 layers, and use 1000 pairs of demonstrations.
Mistral 7B Model. For the best linear SEA setting, we use 1000 pairs of demonstrations randomly sampled from our training split of BBQ to obtain the editing projections for fairness. We set the hyperparameter $K=99.9\%$ and we edit the top 1 layer. For the best $\Phi-$SEA setting, we use squared-exponential feature function, $K=99.99\%$, edit the top 1 layer, and use 1000 pairs of demonstrations.
D.3 Used Prompt Templates for Varied LLM Families
Given the variety of LLM families in our experiment, each employs distinct prompt templates. We adhere to the provided template for each family, as outlined in Table 5.
Table 5: Prompt templates used in our experiments. {system prompt} refers to the provided system prompt for LLaMA-2-Chat model in [37], whereas {user message} refers to our actual input prompts.
LLM Family Prompt Template
LLaMA [INST]<<SYS>>
{system prompt}
<</SYS>>
{user message}[/INST]
Gemma <start_of_turn>user
{user message}<end_of_turn>
<start_of_turn>model
Mistral [INST] {user message} [/INST]
D.4 Experiments Setup for Control Tasks
@@
Appendix E Compute Resources
All experiments for LLaMA-2-Chat-7B are conducted on a single CPU machine (Intel® Xeon® Platinum 8360Y CPUs), utilising 32 cores per experiment, with one 40GB NVIDIA A100 Tensor Core GPU.
Calculation of Editing Projections. When run on a GPU, the computation time for calculating the 21-layer linear SEA for truthfulness is approximately 2 minutes and 32 seconds. The computation time for the 3-layer $\Phi$-SEA for fairness is approximately 20 seconds.
Inference on Benchmarks. For TruthfulQA, the overall inference time of SEA-edited 7B LLaMA-2-Chat model is approximately 10 minutes. For BBQ, the overall inference time for the linear and non-linear SEA is approximately 19 and 21 minutes, respectively.
Using SEA with Other LLMs. Applying SEA to other LLMs incurs minimal additional computational overhead. For the 13B model, SEA implementation necessitates only two A100 40G GPUs. Similarly, for the 70B model, SEA requires just two A100 80G GPUs. These requirements are the same as the usage of LLMs without applying SEA.
Appendix F Limitations
@@
Appendix G Broader Impact
Large language models (LLMs) have had a transformative impact on natural language processing, but their tendency to generate inaccurate or biased content has been a major obstacle to trusted real-world deployment. Our proposed spectral editing of activations (SEA) method aims to improve the truthfulness and fairness of LLMs while maintaining high inference efficiency. If successful, this could help mitigate the spread of misinformation and promote more equitable AI systems.
Improving the factual accuracy of LLM generations has great potential for positive societal impact. Misinformation and disinformation spread via natural language can cause significant harm, eroding trust in institutions, fomenting social divides, and enabling the proliferation of conspiratorial thinking. By making LLMs more truthful, SEA could help curb the flow of inaccurate information from AI systems as they become more widely deployed for language tasks. Additionally, enhancing fairness and reducing encoded biases in LLMs promotes AI that is more inclusive and does not unfairly disadvantage or discriminate against certain groups based on attributes like race or gender.
However, there are also potential negative impacts to consider. Any technology that can steer language model outputs, even with the positive intent of improving truthfulness and fairness, could potentially be misused to amplify or instil other undesirable traits. There are security implications if the editing of LLM activations enables new attack vectors or model vulnerabilities. Care must also be taken with the human demonstrations used to exemplify positive and negative behaviours, as these could perpetuate societal biases present in the data.
While this work is primarily foundational research into a novel model editing technique, we acknowledge the need for proactive consideration of potential negative impacts. As the work progresses towards application, it will be critical to implement robust monitoring, evaluation, and mitigation strategies to uphold principles of responsible AI development and deployment. This includes carefully auditing the data used for supervision, testing outputs across different demographic groups, and implementing appropriate controls against misuse or unintended negative consequences.
Appendix H Demonstrations for TruthfulQA and BBQ